OpenStudy (anonymous):

A circular jogging track is being designed so that one complete lap around the track is one mile long. What should the radius of the track be to the nearest foot? (one mile equals five thousand two hundred eighty feet)

OpenStudy (anonymous):

r = (C/2*pi) r = [5280/(2*3.14)] r = 841 feet